OpenStudy (anonymous):

How to find the vertex of y = x2 − 2x

jimthompson5910 (jim_thompson5910):

First find the x coordinate of the vertex x = -b/(2a) x = -(-2)/(2*1) x = 2/2 x = 1 Then plug this value into y = x^2 - 2x to find the y coordinate of the vertex.